OpenStudy (lexy148):

how many four digit whole numbers greater than 5,000 can be formed using each of the digits 3,5,6,8 once in each number

OpenStudy (anonymous):

three choices for the first number (since it can't be 1) then 3 for the second, two for the third and one for the last by the counting principle, multiply these together \[3\times 3\times 2\times 1\]

OpenStudy (lexy148):

it would be 18